# Global Behavior of Learning Dynamics in Zero-Sum Games with Memory Asymmetry

### v1

@@ 5.2 Global convergence in more advanced zero-sum games
First, Fig. 4A shows the case where the other elements of the payoff matrix are random numbers following the uniform distribution of $[-1,1]$. In this case, the payoff matrix does not degenerate. In other words, X has no two different strategies that generate equal payoffs for all Y’s strategies. In mathematics, there exists no $\bm{a}=(a_{i})_{1\leq i\leq m_{\sf X}}\in\mathbb{R}^{m_{\sf X}}$ such that $\Sigma_{i}a_{i}\bm{u}_{i}=\bm{0}$. We observe that the single Nash equilibrium exists, and Y’s strategy always converges to its equilibrium independent of X’s and Y’s initial strategies.
@@
Second, Fig. 4B shows the case where all the other elements take 00. In this case, the payoff matrix degenerates; Indeed, $\bm{u}_{i}+\bm{u}_{i^{\prime}}=\bm{0}$ holds for $(i,i^{\prime})=(1,3)$, $(2,4)$. Continuous Nash equilibria are seen there, $\bm{x}^{*}=r_{\sf X}(0,1/2,0,1/2)+(1-r_{\sf X})(1/2,0,1/2,0)$ and $\bm{y}^{*}=r_{\sf Y}(0,1/2,0,1/2)+(1-r_{\sf Y})(1/2,0,1/2,0)$ for all $0\leq r_{\sf X}$ and $0\leq r_{\sf Y}\leq 1$. We observe that Y’s strategy converges to one of its equilibria depending on X’s and Y’s initial strategies.
@@
Third, Fig. 4C shows the case where the other elements take 00 in principle except for $u_{11}=-1$. In this case, the payoff matrix degenerates because $\bm{u}_{i}+\bm{u}_{i^{\prime}}=\bm{0}$ holds for $(i,i^{\prime})=(2,4)$. Furthermore, the only Nash equilibrium exists on the boundary of strategy spaces, $\bm{x}^{*}=(0,1/2,0,1/2)$, and $\bm{y}^{*}=(1/2,0,1/2,0)$. As far as we experiment, when the payoff matrix degenerates and only the boundary Nash equilibrium exists, Y’s strategy does not converge to its equilibrium exceptionally. Nevertheless, we observe that X’s strategy converges to its equilibrium instead of Y’s.
6 Conclusion
@@
This study considered the simplest situation of memory asymmetry between two players; only player X memorizes the other’s previous action, while player Y cannot. We formulated their learning dynamics based on the replicator dynamics. Although the existence of memory complicates the dynamics, we captured the global behavior of the learning dynamics by introducing two new quantities. One is the conditional-sum divergence, which is an extension of the previous divergence to the case of games of a reactive strategy. The time change of this extended divergence is characterized by the definiteness of $\bm{X}^{\rm T}\bm{U}$. The other is a family of Lyapunov functions, characterized by $\bm{U}\log\bm{X}^{\rm T}$. This is proved to monotonically increase with time. As a valid application of these two quantities, we proved the global convergence in the learning dynamics in the simplest game equipped with an interior Nash equilibrium. We further experimentally observed the global convergence in more advanced zero-sum games. It is still a conjecture whether the learning dynamics in games of a reactive strategy converge to the Nash equilibrium even in general zero-sum payoff matrices. This study provides novel and valid indicators to analyze dynamics in learning in games.
@@ Appendix A Proofs
$=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}\left(\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}-\Sigma_{i}x_{i|j}\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}\right)-\Sigma_{j}y_{j}^{*}\left(\frac{\partial u^{{\rm st}}}{\partial y_{j}}-\Sigma_{j}y_{j}\frac{\partial u^{{\rm st}}}{\partial y_{j}}\right)$
@@
$=$ $-\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i}^{*}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(A)}=u^{*}}+\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i|j}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(B)}=u^{{\rm st}}}+\underbrace{\Sigma_{i}\Sigma_{j}y_{j}^{*}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{*}}-\underbrace{\Sigma_{i}\Sigma_{j}y_{j}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{{\rm st}}}$
@@
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\left(\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}-\Sigma_{i}x_{i|j}\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}\right)\delta_{i^{\prime}}$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}(\Sigma_{j^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}-\Sigma_{i^{\prime\prime}}x_{i^{\prime\prime}|j}\Sigma_{j^{\prime}}u_{i^{\prime\prime}j^{\prime}}y_{j^{\prime}})$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}-\Sigma_{i}\Sigma_{i^{\prime\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\underbrace{(\Sigma_{i^{\prime}}\delta_{i^{\prime}})}_{=0}\delta_{i}u_{ij}y_{j}x_{i^{\prime\prime}|j}u_{i^{\prime\prime}j^{\prime}}y_{j^{\prime}}$
@@
This equation shows that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ always holds. Furthermore, $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$ is satisfied if and only if ${\rm d}\bm{y}=\bm{0}\Leftrightarrow\bm{y}=\bm{y}^{*}$. ∎
A.3 Proof of Thm. 3
@@ Appendix C Connection to the classical divergence
Appendix D Computational environment
The simulations presented in this paper were conducted using the following computational environment.
  • Operating System: macOS Monterey (version 12.4)
  • Programming Language: Python 3.11.3
  • Processor: Apple M1 Pro (10 cores)
  • Memory: 32 GB

### v2

@@ 4.2. Example: Coupled Matching Pennies
First, Fig. 4A shows the case where the other elements of the payoff matrix are random numbers following the uniform distribution of $[-1,1]$. In this case, the payoff matrix is linearly independent. In mathematics, there exists no $\bm{a}=(a_{j})_{1\leq j\leq m_{\sf Y}}\in\mathbb{R}^{m_{\sf Y}}$ such that $\Sigma_{j}a_{j}\bm{u}_{j}=\bm{0}$ other than $\bm{a}=\bm{0}$. Thus, there is a single Nash equilibrium in the interior of the strategy space. As in the matching pennies game, we observe that Y’s strategy always converges to its equilibrium independent of X’s and Y’s initial strategies.
Continuous equilibrium:
@@
Second, Fig. 4B shows the case where all the other elements take 00. In this case, the payoff matrix is not linearly independent in two ways. Indeed, $\Sigma_{j}a_{j}\bm{u}_{j}=\bm{0}$ for $\bm{a}=(0,1,0,1)$ and $(1,0,1,0)$. Thus, Nash equilibria exist continuously as $\bm{x}^{*}=r_{\sf X}(0,1/2,0,1/2)+(1-r_{\sf X})(1/2,0,1/2,0)$ and $\bm{y}^{*}=r_{\sf Y}(0,1/2,0,1/2)+(1-r_{\sf Y})(1/2,0,1/2,0)$ for all $0\leq r_{\sf X}\leq 1$ and $0\leq r_{\sf Y}\leq 1$. Even in such continuous equilibria, we observe that Y’s strategy converges to one of the equilibria depending on X’s and Y’s initial strategies.
Boundary equilibrium:
@@
Third, Fig. 4C shows the case where the other elements take 00 in principle except for $u_{11}=-1$. In this case, the payoff matrix is not linearly independent in one way. Indeed, $\Sigma_{j}a_{j}\bm{u}_{j}=\bm{0}$ for $\bm{a}=(0,1,0,1)$. The only Nash equilibrium exists on the boundary of strategy spaces, $\bm{x}^{*}=(0,1/2,0,1/2)$, and $\bm{y}^{*}=(1/2,0,1/2,0)$. As far as we observe our experiments, Y’s strategy fails to converge the equilibrium when the payoff matrix is not linearly independent and is equipped with only the boundary Nash equilibrium. Nevertheless, we observe that X’s strategy converges to the equilibrium instead of Y’s. We find no case that neither X’s nor Y’s strategy converges to the equilibrium.
5. Conclusion
@@
This study considered the simplest situation of memory asymmetry between two players; only player X memorizes the other’s previous action, while player Y cannot. We formulated their learning dynamics based on the replicator dynamics. Although the existence of memory complicates the dynamics, we captured the global behavior of the learning dynamics by introducing two new quantities. One is the conditional-sum divergence, which is an extension of the previous divergence to the case of reactive strategies. We proved that when X exploits Y, this conditional-sum divergence becomes smaller, meaning that their strategies converge to the Nash equilibrium. The other is a family of Lyapunov functions, meaning X’s exploitability to Y. We proved that these Lyapunov functions monotonically increase, meaning that X learns to exploit Y with time. As a valid application of the combination of these two quantities, we suggested the global convergence to the Nash equilibrium. Theoretically, we proved the global convergence in the matching pennies game, the simplest game equipped with an interior Nash equilibrium. Our experiments further support that global convergence occurs in coupled matching pennies games, which can have various types of Nash equilibrium structures, such as interior equilibrium, continuous equilibrium, and boundary equilibrium. It is still a conjecture whether the learning dynamics with memory asymmetry converge to the Nash equilibrium in general zero-sum games. This study provides novel and valid indicators to analyze dynamics in learning in games with memories.
{acks}
K. Ariu is supported by JSPS KAKENHI Grant No. 23K19986.
@@ Appendix A Proofs
$=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}x_{i|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right)-\Sigma_{j}y_{j}^{*}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}y_{j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right)$
@@
$=$ $-\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i}^{*}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(A)}=u^{*}}+\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i|j}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(B)}=u^{{\rm st}}}+\underbrace{\Sigma_{i}\Sigma_{j}y_{j}^{*}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{*}}-\underbrace{\Sigma_{i}\Sigma_{j}y_{j}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(D)}=u^{{\rm st}}}$
@@
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}-\Sigma_{i^{\prime\prime}}x_{i^{\prime\prime}|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\right)\delta_{i^{\prime}}$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}-\Sigma_{i}\Sigma_{j}\delta_{i}u_{ij}\Sigma_{i^{\prime\prime}}x_{i^{\prime\prime}|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\underbrace{\Sigma_{i^{\prime}}\delta_{i^{\prime}}}_{=0}$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}$
@@
This equation shows that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ always holds.
Furthermore, let us prove that $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$ is equivalent to ${\rm d}\bm{y}=\bm{0}\Leftrightarrow\bm{y}=\bm{y}^{*}$. Here, we substitute $\bm{\delta}=\bm{x}-\bm{x}^{*}$ for $\bm{x}\in\Delta^{m_{\sf X}-1}$ (but no problem occurs even if we generally consider $\bm{x}\in\mathbb{R}^{m_{\sf X}}$ such that $\Sigma_{i}x_{i}=1$), and then
$\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y}$ $=(\bm{x}-\bm{x}^{*})^{\rm T}\bm{U}(\bm{y}-\bm{y}^{*})$ (25)
$=\underbrace{\bm{x}^{\rm T}\bm{U}\bm{y}}_{=u^{{\rm st}}(\bm{x},\bm{y})}-\underbrace{\bm{x}^{\rm T}\bm{U}\bm{y}^{*}}_{=u^{*}}-\underbrace{\bm{x}^{*\rm T}\bm{U}\bm{y}}_{=u^{*}}+\underbrace{\bm{x}^{*\rm T}\bm{U}\bm{y}^{*}}_{=u^{*}}$ (26)
$=u^{{\rm st}}(\bm{x},\bm{y})-u^{*}.$ (27)
We obtain
$H^{\dagger}(\bm{y};\bm{x}-\bm{x}^{*})=|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|^{2}.$ (28)
We use the proof by contradiction. Assume ${\rm d}\hat{\bm{y}}\neq\bm{0}\Leftrightarrow\hat{\bm{y}}\neq\bm{y}^{*}$ such that $u^{{\rm st}}(\bm{x},\hat{\bm{y}})=u^{*}$ for all $\bm{x}$, and then we obtain for $\bm{x}=\bm{x}^{*}$
$u^{{\rm st}}(\bm{x}^{*},\hat{\bm{y}})=u^{*}=\max_{\bm{y}}u^{{\rm st}}(\bm{x}^{*},\bm{y})$ (29)
$\Leftrightarrow\hat{\bm{y}}=\arg\max_{\bm{y}}u^{{\rm st}}(\bm{x}^{*},\bm{y})=\bm{y}^{*},$ (30)
which contradicts $\hat{\bm{y}}\neq\bm{y}^{*}$. We proved $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$ is equivalent to ${\rm d}\bm{y}=\bm{0}\Leftrightarrow\bm{y}=\bm{y}^{*}$. ∎
A.3. Proof of Cor. 4
@@ Appendix C Connection to the classical divergence
Appendix D Extension to other learning algorithms
In the main text, we focus only on the replicator dynamics. However, our findings are also applicable to the gradient descent-ascent algorithm. In the interior of the strategy space, the learning dynamics of the gradient descent-ascent are described as
$\dot{x}_{i|j}=+\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right),\quad\dot{y}_{j}=-\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}\frac{1}{m_{\sf Y}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right),$ (41)
which correspond to Eqs. (5) and (6), respectively.
To capture the global behavior in the learning dynamics based on the gradient descent-ascent, we have to make small changes in $D(\bm{X},\bm{y})$ and $H(\bm{X};\bm{\delta})$ as
$D(\bm{X},\bm{y})$ $:=\Sigma_{j}D_{2}(\bm{x}^{*},\bm{x}_{j})+D_{2}(\bm{y}^{*},\bm{y}),\quad D_{2}(\bm{p}^{*},\bm{p}):=\frac{1}{2}\|\bm{p}-\bm{p}^{*}\|^{2},$ (42)
$H(\bm{X};\bm{\delta})$ $:=\bm{\delta}^{\rm T}\bm{U}\bm{X}^{\rm T}\bm{\delta},$ (43)
which correspond to Eqs. (13) and (15), respectively.
This $D(\bm{X},\bm{y})$ satisfies Thm. 2 as follows.
$\dot{D}(\bm{X},\bm{y})$ $=\Sigma_{i}\Sigma_{j}\dot{x}_{i|j}(x_{i|j}-x_{i}^{*})+\Sigma_{j}\dot{y}_{j}(y_{j}-y_{j}^{*})$
$=\Sigma_{i}\Sigma_{j}(x_{i|j}-x_{i}^{*})\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}+-\Sigma_{i}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right)-\Sigma_{j}(y_{j}-y_{j}^{*})\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}\frac{1}{m_{\sf Y}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right)$
$=\Sigma_{i}\Sigma_{j}(x_{i|j}-x_{i}^{*})\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{j}(y_{j}-y_{j}^{*})\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}$
$=\underbrace{\Sigma_{i}\Sigma_{j}(x_{i|j}-x_{i}^{*})y_{j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}}_{=({\rm B})-({\rm A})=u^{{\rm st}}-u^{*}}-\underbrace{\Sigma_{j}(y_{j}-y_{j}^{*})\Sigma_{i}u_{ij}x_{i}^{{\rm st}}}_{=:({\rm D})-({\rm C})=u^{{\rm st}}-u^{*}}-\underbrace{\Sigma_{j}(y_{j}-y_{j}^{*})\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}}_{=({\rm E})={\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}}$
$=-{\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$
$=D^{\dagger}(\bm{X};{\rm d}\bm{y}).$ (44)
Here, we used the terms $({\rm A})$-$({\rm E})$ in Appendix A.1.
Furthermore, this $H(\bm{X};\bm{\delta})$ also satisfies Thm. 3 as follows.
$\dot{H}(\bm{X};\bm{\delta})$ $=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\dot{x}_{i^{\prime}|j}\delta_{i^{\prime}}$
$=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}+-\Sigma_{i^{\prime\prime}}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\right)\delta_{i^{\prime}}$
$=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}+\Sigma_{i}\Sigma_{j}\delta_{i}u_{ij}\Sigma_{i^{\prime\prime}}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\underbrace{\Sigma_{i^{\prime}}\delta_{i^{\prime}}}_{=0}$
$=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}$
$=H^{\dagger}(\bm{y};\bm{\delta}).$ (45)
In the final line, we used the same calculation as that of Appendix A.2.
Both the replicator dynamics and the gradient descent-ascent are included in the Follow the Regularized Leaders. A promising future study is to capture the global behavior in learning dynamics based on general Follow the Regularized Leaders algorithms with memories.